# Can Large Language Models Beat Wall Street? Unveiling the Potential of AI in Stock Selection

### v1

@@ 3.1 Progressive News Summarizer
The Progressive News Summarizer is responsible for news acquisition, condensation, and crafting a progressive synopsis of a stock’s most influential news. As portrayed in Figure 2, daily news items corresponding to a specific stock are fetched from available APIs. The study utilized EODHD [^2] https://eodhd.com/financial-apis/stock-market-financial-news-api/ Stock Market and Financial News API for news sourcing.
The daily news for a company is preprocessed to exclude text unrelated to the company, such as clickbait articles, ensuring it is in an appropriate format for entry into the prompt. GPT-4, accessed via OpenAI’s API, is systematically prompted to distill the stock’s daily news and generate a concise daily news summary ($N_{S,day}$). The latter is stored in a centralized repository.
@@
While this method provides a summary for a specific date, it is essential to include the ongoing narrative of news-related content for the company, particularly older news that remains significant. For example, in the case of a merger or legal dispute, an announcement about the company’s better-than-expected results may be less influential in the overall decision-making process. The Progressive News Summarizer addresses this by integrating the latest news summaries ($N_{S,days-t}$) with the preceding progressive summary ($PN_{S,t-1}$). The GPT-4 model generates a curated narrative that incorporates key events, becoming instrumental in MarketSenseAI’s subsequent analyses.
In this work, the Progressive News Summarizer is designed to operate on a monthly basis, ensuring that it captures the most current and pertinent state of a company as reflected in news coverage. Specifically, on the last day of each month, following market closure, it generates an updated summary for each stock. This summary is a sophisticated synthesis that takes into account both the daily news summaries of the past month and the Progressive News Summary formulated in the previous month. This process ensures that the summarizer consistently reflects the latest, most relevant, and significant developments, offering a comprehensive and current snapshot of the company’s status in the news.
Moreover, the flexibility of the Progressive News Summarizer’s design allows for its frequency to be modified to suit diverse investment strategies. It can be configured for different intervals, such as biweekly, weekly, or even daily, thereby accommodating various temporal requirements and investment strategies.
@@
In October, the focus was on Apple’s significant role in the tech industry, marked by the launch of the iPhone 15 series and updates to its Apple Watch and AirPods. This period also highlighted challenges in smartphone sales, the impact of geopolitical issues, and fluctuations in stock performance. Unique to this month were reports on Apple’s interest in acquiring Formula 1 broadcasting rights and CEO Tim Cook’s sale of company stock, underscoring the company’s diverse strategic interests and executive decisions.
@@
Financial Performance
Yes
Yes
Guidance for Q4 2023
Yes
Yes
iPhone 15 Launch
Yes
Yes
New Apple Watch and AirPods
Yes
No
Smartphone Sales Challenges
Yes
Yes
Geopolitical Issues Impact
Yes
Yes
Stock Performance
Yes
Yes
Partnership with DuckDuckGo
Yes
No
CEO Tim Cook’s Stock Sale
Yes
No
Interest in Formula 1 Broadcasting
Yes
No
Product Lineup including M3 Chip
No
Yes
Sales Slowdown and Competition
No
Yes
Appeal Win in UK
No
Yes
Partnership with Goldman Sachs
No
Yes
Regulatory Issues in Payment Apps
No
Yes
@@
Market Position
Dominant in the tech sector with record services revenue and robust product lineup, including the iPhone 15 series and new Mac products.
Sales Performance
Experiencing a slowdown, potentially dropping 5% in iPhone sales due to challenges in China and Japan.
Stock Analysis
Recent dip viewed as a buying opportunity by analysts, seasonal performance aligns with product launch cycle.
Legal and Strategic Moves
Won UK appeal on mobile browser and cloud gaming services, potentially ending credit card partnership with Goldman Sachs.
Innovation and Sustainability
Launched M3 silicon chip based on 3-nanometer technology, expanding streaming content on Apple TV+.
Regulatory Challenges
Faces oversight of digital wallets and payment apps, navigating geopolitical and economic risks.
Investment Consideration
Despite challenges, presents a promising opportunity with strategic expansion, innovative products, and strong services division.
@@ 3.2 Fundamentals Summarizer
Fundamental data is crucial in predictive financial analytics, offering quantifiable metrics that reflect a company’s current health and future trajectory. As depicted in Figure 3, we utilize EODHD’s Fundamental Data API to source this quarterly information. To facilitate the comparison of financial data, we preprocess the data before inputting it into the prompt. Specifically, we employ numerical abbreviation techniques, using prefixes like ”million” and ”billion” to represent large numbers in a more compact and readable format. Additionally, data from different quarters are placed side by side in a table format. The resulting prompt, fed to the GPT-4 model, delves into aspects such as profitability, revenue trajectory, debt metrics, and cash flow dynamics by comparing the two most recent quarterly financial statements. This focus on recent data enables the LLM to detect shifts in financial performance, potentially correlating it with progressive news. Although LLMs traditionally face challenges in comprehending complex numerical data, our preprocessing approach, combined with GPT-4’s capabilities, ensures accurate comparison and interpretation. The fundamentals summarizer is designed to present an unbiased, factual overview of a company’s financial status, avoiding any direct investment recommendation.
@@
Table 3 showcases how the Fundamentals Summarizer distills key insights from financial statements, including income, balance, and cash flow statements. Similar to the Progressive News Summarizer, this component can be employed independently to provide a concise financial overview of the company being analyzed.
@@
Profitability
Apple’s net income has increased from \$19.88 billion in Q2 2023 to \$22.96 billion in Q3 2023, indicating a strong profitability. The company’s gross profit also increased from \$36.41 billion to \$40.43 billion over the same period.
Revenue Growth
Total revenue increased from \$81.80 billion in Q2 2023 to \$89.50 billion in Q3 2023, showing a positive revenue growth
Debt Levels
The company’s total liabilities increased from \$274.76 billion in Q2 2023 to \$290.44 billion in Q3 2023. The long-term debt decreased slightly from \$98.07 billion to \$95.28 billion, but the short-term debt increased from \$11.21 billion to \$15.81 billion. The net debt also increased slightly from \$80.87 billion to \$81.12 billion. This indicates that the company’s debt level is increasing, which could be a concern if not managed properly.
Cash Flow
The company’s net cash from operating activities decreased from \$26.38 billion in Q2 2023 to \$21.60 billion in Q3 2023. However, the end period cash flow increased from \$29.90 billion to \$30.74 billion, indicating a positive cash flow generation.
Assets and Equity
Total assets increased from \$335.04 billion in Q2 2023 to \$352.58 billion in Q3 2023. The total stockholder equity also increased from \$60.27 billion to \$62.15 billion, indicating a growth in the company’s assets and equity.
Conclusion
Apple Inc. shows strong profitability and revenue growth. However, the increasing debt level needs to be monitored. The company has a positive cash flow generation, and its assets and equity are growing.
@@ 3.3 Stock Price Dynamics Summarizer
The Stock Price Dynamics Summarizer, a key component of MarketSenseAI, analyzes and contextualizes the price movements and financial metrics of stocks. As shown in Figure 4, this component not only examines the target stock but also compares its performance with the five most similar stocks, based on company description and sector, and includes the broader market context represented by the S&P 500 index.
@@
The methodology for identifying similar stocks is outlined in Algorithm 1, utilizing the MPNet language model to generate embeddings and compute similarity scores (Song et al., 2020). This approach ensures a thorough and relevant comparative analysis, integrating both individual stock performance and broader market trends.
@@
Cumulative Return
Apple Inc demonstrated a 29.0% return, outperforming the S&P 500 index’s 13.7% but underperforming tech peers like Adobe Systems and Amazon.com Inc.
Sharpe Ratio
Apple’s Sharpe Ratio of 1.34 indicates a favorable risk-adjusted return compared to the market index Sharpe Ratio of 0.99, suggesting better compensation for the risk taken.
Volatility
Apple’s volatility at 21.7% is lower than that of Alphabet Inc, Adobe, and Amazon, indicating less erratic stock price movements.
Maximum Drawdown
Apple experienced a maximum drawdown of -16.0%, which is less severe than the drawdowns of Adobe, Amazon, and Best Buy Co. Inc.
Correlation
Apple shows a high correlation with the S&P 500 (0.76) and moderate correlation with other tech stocks like Microsoft Corporation and Amazon.
Conclusion
Apple has shown resilience and strong risk-adjusted performance relative to the broader market and some tech peers, with lower volatility and a relatively modest maximum drawdown.
@@ 3.4 Macroeconomic Environment Summary
Conducting an in-depth macroeconomic analysis is essential for making informed investment decisions and effective capital allocation. Such analysis offers vital insights into the overall economic health and performance, significantly influencing the profitability and value of individual companies as well as the broader stock market. By considering major forces that shape the investment landscape, such as the Covid-19 pandemic or the Ukraine war, investors can make better-informed decisions.
To facilitate this, MarketSenseAI includes a component named MarketDigest, depicted in Figure 5. This component synthesizes investment reports and research articles biweekly, providing succinct summaries of complex economic data and trends. MarketDigest sources information from a variety of publicly accessible reports from leading banks and investment institutions, including Goldman Sachs, Morgan Stanley, UBS, and BlackRock.
@@
The process begins by transforming these reports and articles into text form. Subsequently, the component utilizes GPT-4 to summarize individual reports, and then, in a second step, it condenses these individual summaries into a comprehensive overview. This approach allows MarketDigest to encapsulate diverse perspectives and analyses into a coherent narrative, offering a consensus view on the macroeconomic climate, central bank policies, preferred sectors or countries, as well as geopolitical trends. The output is concise yet thorough, accounting for potential contradictions or differing viewpoints among market analysts and experts.
Table 5 provides an example of a MarketDigest summary. Notably, KM Cube Asset Management [^3] https://www.km3am.com/ already disseminates MarketDigest as a bi-weekly newsletter to its clientele, underscoring its practical application and value in the real-world financial context (Metaxas, 2023).
@@
Inflation
US core PCE inflation eased to 3.5% in October, indicating a disinflation trend.
Interest Rates
Market-implied pricing suggests potential rate cuts in March 2024 for both the US and Europe.
Japan’s Monetary Policy
Bank of Japan expected to weaken or abandon yield curve control due to domestic inflation.
ECB Policy
European Central Bank has begun balance sheet unwind.
Bond Market Outlook
Positive outlook on short- to medium-term developed market sovereign bonds.
Equity Market Stance
Neutral stance on developed market equities, with US stocks as largest allocation.
Global Growth
Global economy expected to experience below-trend growth in 2024.
Investment Strategy
Portfolios should maintain neutral exposure to risk and equities, overweight allocation to quality fixed income.
US Dollar
US dollar’s position as leading global reserve currency showing signs of vulnerability.
Employment Risks
Risks to employment are on the downside, with leading indicators of employment deteriorating significantly.
Market Rally
Global financial markets experiencing significant rally, boosted by cooling inflation and falling Treasury yields.
Contradictions
Positive outlook on bonds but neutral on equities; US dollar vulnerability but remains a key currency.
Positive Sentiment
Short- to medium-term bonds, inflation-linked bonds, private market income, quality fixed income, US stocks.
Negative Sentiment
Credit, US Treasury, private markets, small-cap equities, Chinese equities.
Neutral Sentiment
Developed market equities, investment-grade credit, real estate, private equity funds, emerging markets outside China.

### v2

@@ 3.1 Progressive News Summarizer
The Progressive News Summarizer ($PN_{S,t}$) is responsible for news acquisition, condensation, and crafting a progressive synopsis of a stock’s most influential news. As portrayed in Figure 2, daily news items corresponding to a specific stock are fetched from available APIs. The study utilized EODHD [^2] https://eodhd.com/financial-apis/stock-market-financial-news-api/ Stock Market and Financial News API for news sourcing. Equation 1 and 2 model the processes for obtaining the daily and the progressive news summaries, respectively.
$N_{S,\tau}=\bigoplus_{i=t-\tau}^{t}\text{Summarize}(N_{S,i})$ (1)
$PN_{S,t}=\text{Summarize}(PN_{S,t=t-1},N_{S,\tau})$ (2)
Where,
[…]
representing the time window for news
progressive summary.
[…]
daily news summaries.
The daily news for a company is preprocessed to exclude text unrelated to the company, such as clickbait articles, thus ensuring it is in an appropriate format for entry into the prompt. GPT-4, accessed via OpenAI’s API, is systematically prompted to distill the stock’s daily news and generate a concise daily news summary ($N_{S,i}$). The latter is stored in a centralized repository.
@@
While this method provides a summary for a specific date, it is essential to include the ongoing narrative of news-related content for the company, particularly older news that remains significant. For example, in the case of a merger or legal dispute, an announcement about the company’s better-than-expected results may be less influential in the overall decision-making process. The Progressive News Summarizer addresses this by integrating the latest news summaries ($N_{S,\tau}$) with the preceding progressive summary ($PN_{S,t-1}$). More specifically, the prompt structure instructs GPT-4 to mimic the role of a financial analyst with the task of synthesizing an updated summary for a specific stock by integrating various types of information. The prompt includes:
  • 1.
    Current Summary: The latest summary of the company and its stock as of a specific month and year.
  • 2.
    Daily News Summary: News articles, divided into factual news and analysts’ opinions for the company over a given month.
  • 3.
    Instructions: Integrate the most pertinent information, distinguish factual news and analysts’ opinions.
This summary is a synthesis that takes into account both the daily news summaries of the past month and the Progressive News Summary formulated in the previous month. This process ensures that the summarizer consistently reflects the latest, most relevant, and significant developments, offering a comprehensive and current snapshot of the company’s status in the news.
While the Progressive News Summarizer in this study is tailored for monthly intervals, its design offers adaptability for various frequencies to match different investment approaches. The effects of adjusting the summarizer’s frequency for different intervals on the effectiveness of diverse investment strategies, remain an area for future exploration.
@@
In October, the focus was on Apple’s significant role in the tech industry, marked by the launch of the iPhone 15 series and updates to its Apple Watch and AirPods. This period also highlighted challenges in smartphone sales, the impact of geopolitical issues, and fluctuations in stock performance. Unique to this month were reports on Apple’s interest in acquiring Formula 1 broadcasting rights and the sale of company stock by CEO Tim Cook, underscoring the company’s diverse strategic interests and executive decisions.
@@
Financial Performance Yes Yes
Guidance for Q4 2023 Yes Yes
iPhone 15 Launch Yes Yes
New Apple Watch and AirPods Yes No
Smartphone Sales Challenges Yes Yes
Geopolitical Issues Impact Yes Yes
Stock Performance Yes Yes
Partnership with DuckDuckGo Yes No
CEO Tim Cook’s Stock Sale Yes No
Interest in Formula 1 Broadcasting Yes No
Product Lineup including M3 Chip No Yes
Sales Slowdown and Competition No Yes
Appeal Win in UK No Yes
Partnership with Goldman Sachs No Yes
Regulatory Issues in Payment Apps No Yes
@@
Market Position Dominant in the tech sector with record services revenue and robust product lineup, including the iPhone 15 series and new Mac products.
Sales Performance Experiencing a slowdown, potentially dropping 5% in iPhone sales due to challenges in China and Japan.
Stock Analysis Recent dip viewed as a buying opportunity by analysts, seasonal performance aligns with product launch cycle.
Legal and Strategic Moves Won UK appeal on mobile browser and cloud gaming services, potentially ending credit card partnership with Goldman Sachs.
Innovation and Sustainability Launched M3 silicon chip based on 3-nanometer technology, expanding streaming content on Apple TV+.
Regulatory Challenges Faces oversight of digital wallets and payment apps, navigating geopolitical and economic risks.
Investment Consideration Despite challenges, presents a promising opportunity with strategic expansion, innovative products, and strong services division.
@@ 3.2 Fundamentals Summarizer
Fundamental data is crucial in predictive financial analytics, offering quantifiable metrics that reflect a company’s current health and future trajectory. As depicted in Figure 3, we utilize EODHD’s Fundamental Data API to source this quarterly information. To facilitate the comparison of financial data, we preprocess the data before inputting it into the prompt. This preprocessing includes a specific numerical abbreviation technique, which converts large numbers into a more compact format by representing them with prefixes like ”million”, ”billion”, or ”thousand”. For example, numbers in the billions are formatted as ’X billion’, where X is the original number divided by 1 billion, rounded to two decimal places. This approach standardizes the financial data, ensuring consistency and clarity in the inputs fed to the LLM. We found that this preprocessing step is crucial for allowing GPT-4 to accurately compare and interpret complex financial figures.
Additionally, data from different quarters are placed side by side in a table format. The resulting prompt, fed to the GPT-4 model, delves into aspects such as profitability, revenue trajectory, debt metrics, and cash flow dynamics by comparing the most recent quarterly financial statements. This focus on recent data enables the LLM to detect shifts in financial performance, potentially correlating it with progressive news. More specifically, MarketSenseAI models the financial condition of company using the following formula:
$F_{s}=\text{Summarize}\left(\text{Standardize}\left(\bigcup_{i=1}^{n}\text{FinancialData}_{s,q_{i}}\right)\right)$ (3)
Where,
[…]
where $\quad\text{where }i=1,\dots,n\text{ (the last }n\text{ quarters).}$).
[…]
numerical abbreviation.
[…]
from this standardized data.
The prompt structure guides GPT-4 to adopt the role of a financial analyst focusing on recent trends. The AI is tasked with evaluating the financial health of a specified company’s stock by analyzing its latest quarters. The prompt includes:
  • 1.
    Financial Tables: Key financial data from Balance Sheet, Income Statement, and Cash Flow for the latest quarters.
  • 2.
    Analysis Focus: Recent trends and developments in profitability, revenue growth, debt levels, and cash flow generation.
  • 3.
    Instructions: Conduct a bullet-form analysis.
Although LLMs traditionally face challenges in comprehending complex numerical data, our preprocessing approach, combined with GPT-4’s capabilities, ensures accurate comparison and interpretation. The fundamentals summarizer is designed to present an unbiased, factual overview of a company’s financial status, avoiding any direct investment recommendation. Table 3 showcases how the Fundamentals Summarizer distills key insights from financial statements, including income, balance, and cash flow statements. Similar to the Progressive News Summarizer, this component can be employed independently to provide a concise financial overview of the company being analyzed.
@@
Profitability Apple’s net income has increased from \$19.88 billion in Q2 2023 to \$22.96 billion in Q3 2023, indicating a strong profitability. The company’s gross profit also increased from \$36.41 billion to \$40.43 billion over the same period.
Revenue Growth Total revenue increased from \$81.80 billion in Q2 2023 to \$89.50 billion in Q3 2023, showing a positive revenue growth
Debt Levels The company’s total liabilities increased from \$274.76 billion in Q2 2023 to \$290.44 billion in Q3 2023. The long-term debt decreased slightly from \$98.07 billion to \$95.28 billion, but the short-term debt increased from \$11.21 billion to \$15.81 billion. The net debt also increased slightly from \$80.87 billion to \$81.12 billion. This indicates that the company’s debt level is increasing, which could be a concern if not managed properly.
Cash Flow The company’s net cash from operating activities decreased from \$26.38 billion in Q2 2023 to \$21.60 billion in Q3 2023. However, the end period cash flow increased from \$29.90 billion to \$30.74 billion, indicating a positive cash flow generation.
Assets and Equity Total assets increased from \$335.04 billion in Q2 2023 to \$352.58 billion in Q3 2023. The total stockholder equity also increased from \$60.27 billion to \$62.15 billion, indicating a growth in the company’s assets and equity.
Conclusion Apple Inc. shows strong profitability and revenue growth. However, the increasing debt level needs to be monitored. The company has a positive cash flow generation, and its assets and equity are growing.
@@ 3.3 Stock Price Dynamics Summarizer
The Stock Price Dynamics Summarizer, a key component of MarketSenseAI, analyzes and contextualizes the price movements and financial metrics of stocks. As shown in Figure 4, this component not only examines the target stock but also compares its performance with the five most similar stocks, based on company description and sector, and includes the broader market context represented by the S&P 500 index. The mathematical representation of the Stock Price Dynamics Summary is given by Equation 4.
$P_{S,t}=\text{Summarize}\left(\text{Metrics}_{S,t},\bigcup_{j=1}^{n}\text{Metrics}_{P_{j},t}\right)$ (4)
Where,
[…]
for stock $\quad\text{for stock }S\text{ at time }t.$
[…]
comparison.
[…]
into a concise, factual report.
The prompt structure includes:
  • 1.
    Performance Metrics: Analysis of the stock’s performance using metrics such as Cumulative Returns, Volatility, Sharpe Ratio, Maximum Drawdown, and a Correlation Matrix.
  • 2.
    Comparative Analysis: Comparing the stock’s performance with related stocks and the S&P 500 index.
  • 3.
    Instructions: Summarize the findings in a concise and factual report.
@@
The methodology for identifying similar stocks is outlined in Algorithm 1, utilizing the MPNet language model to generate embeddings and compute similarity scores (Song et al., 2020). More specifically, stock descriptions are encoded by MPNet into high-dimensional vectors, capturing each company’s unique characteristics and activities. This facilitates the computation of pairwise similarity scores among companies listed in the S&P 500. These scores are crucial in identifying stocks with attributes similar to the target stock, thereby ensuring a comprehensive comparative analysis that integrates individual stock performance with broader market trends.
@@
Cumulative Return Apple Inc demonstrated a 29.0% return, outperforming the S&P 500 index’s 13.7% but underperforming tech peers like Adobe Systems and Amazon.com Inc.
Sharpe Ratio Apple’s Sharpe Ratio of 1.34 indicates a favorable risk-adjusted return compared to the market index Sharpe Ratio of 0.99, suggesting better compensation for the risk taken.
Volatility Apple’s volatility at 21.7% is lower than that of Alphabet Inc, Adobe, and Amazon, indicating less erratic stock price movements.
Maximum Drawdown Apple experienced a maximum drawdown of -16.0%, which is less severe than the drawdowns of Adobe, Amazon, and Best Buy Co. Inc.
Correlation Apple shows a high correlation with the S&P 500 (0.76) and moderate correlation with other tech stocks like Microsoft Corporation and Amazon.
Conclusion Apple has shown resilience and strong risk-adjusted performance relative to the broader market and some tech peers, with lower volatility and a relatively modest maximum drawdown.
@@ 3.4 Macroeconomic Environment Summary
Conducting an in-depth macroeconomic analysis is essential for making informed investment decisions and effective capital allocation. Such analysis offers vital insights into the overall economic health and performance, significantly influencing the profitability and value of individual companies as well as the broader stock market. By considering major forces that shape the investment landscape, such as the Covid-19 pandemic or the war in Ukraine, investors can make better-informed decisions.
To facilitate this, MarketSenseAI includes a component named MarketDigest, depicted in Figure 5. This component synthesizes investment reports and research articles biweekly, providing succinct summaries of complex economic data and trends. MarketDigest sources information from a variety of publicly accessible reports from leading banks and investment institutions, including Goldman Sachs, Morgan Stanley, UBS, and BlackRock. The mathematical representation for MarketSenseAI’s Macroeconomic Environment Summary component, MarketDigest, can be formulated as follows:
$M_{t}=\text{Summarize}\left(\bigcup_{j=1}^{N}\text{Summarize}\left(\text{Report}_{j,t}\right)\right)$ (6)
Where,
$M_{t}$ $:\text{Macroeconomic summary at time }t.$
[…]
concise overview.
[…]
$\text{Report}_{j,t}$ $:\text{Investment report or article }j\text{ at time }t.$
[…]
@@
The process begins by transforming these reports and articles into text form. Subsequently, the component utilizes GPT-4 to summarize individual reports, and then, in a second step, it condenses these individual summaries into a comprehensive overview. This approach allows MarketDigest to encapsulate diverse perspectives and analyses into a coherent narrative, offering a consensus view on the macroeconomic climate, central bank policies, preferred sectors or countries, as well as geopolitical trends. The output is concise yet thorough, accounting for potential contradictions or differing viewpoints among market analysts and experts. The prompt structure for MarketDigest includes:
  • 1.
    Initial Summary Focus: Summarization of individual reports, emphasizing critical macroeconomic elements including central bank policies, geopolitical insights, and market outlooks.
  • 2.
    Synthesis and Sentiment Analysis: An in-depth analysis of all reports to extract consensus and divergent views, with focus on sentiment analysis categorized by asset class or investment dimension.
  • 3.
    Instructions: Demanding a detailed and factual report, with emphasis on prevailing market sentiments and analytical categorization by asset class.
Table 5 provides an example of a MarketDigest summary. Notably, KM Cube Asset Management [^3] https://www.km3am.com/ initiated MarketDigest in March 2023 serving as a critical analytical tool, providing succinct market overviews to its clients. Distributed bi-weekly, it facilitates an enhanced understanding of market dynamics for both private and institutional clients, thereby informing their investment decision-making processes (Metaxas, 2023). Furthermore, MarketDigest is integral to the company’s monthly investment committee meetings, where it substantially contributes to the deliberation and formulation of strategies for investment portfolios and the management of discretionary products.
@@
Inflation US core PCE inflation eased to 3.5% in October, indicating a disinflation trend.
Interest Rates Market-implied pricing suggests potential rate cuts in March 2024 for both the US and Europe.
Japan’s Monetary Policy Bank of Japan expected to weaken or abandon yield curve control due to domestic inflation.
ECB Policy European Central Bank has begun balance sheet unwind.
Bond Market Outlook Positive outlook on short- to medium-term developed market sovereign bonds.
Equity Market Stance Neutral stance on developed market equities, with US stocks as largest allocation.
Global Growth Global economy expected to experience below-trend growth in 2024.
Investment Strategy Portfolios should maintain neutral exposure to risk and equities, overweight allocation to quality fixed income.
US Dollar US dollar’s position as leading global reserve currency showing signs of vulnerability.
Employment Risks Risks to employment are on the downside, with leading indicators of employment deteriorating significantly.
Market Rally Global financial markets experiencing significant rally, boosted by cooling inflation and falling Treasury yields.
Contradictions Positive outlook on bonds but neutral on equities; US dollar vulnerability but remains a key currency.
Positive Sentiment Short- to medium-term bonds, inflation-linked bonds, private market income, quality fixed income, US stocks.
Negative Sentiment Credit, US Treasury, private markets, small-cap equities, Chinese equities.
Neutral Sentiment Developed market equities, investment-grade credit, real estate, private equity funds, emerging markets outside China.